A pizza parlour will deliver a pizza take-away order to the customer if s/he lives within a 5 kilometre radius from the pizza parlour. If the customer lives within this radius, it is found that the time taken to receive the pizza is normally distributed with a mean time of 45 minutes and a standard deviation of 8 minutes. Paul has relocated and now lives within the 5 kilometre radius. He orders a pizza from the pizza parlour. The probability that it takes between 43 and 49 minutes for him to receive his pizza is (correct to 2 decimal places)

 


1
Expert's answer
2022-05-24T12:34:14-0400
"P(43<X<49)=P(Z<\dfrac{49-45}{8})"




"-P(Z\le\dfrac{43-45}{8})=P(Z<0.5)"




"-P(Z\le-0.25)\approx0.69146-0.40129"




"\approx0.29"